### v3

@@ A.1 Proof of Theorem 10
Our proof of Theorem 10 combines two well-established techniques for proving uniform bounds on the supremum of a random process in the statistical learning theory: chaining, which constructs a sequence of progressively finer approximations to the continuous process, and Sauer-Shelah Lemma, a combinatorial principle using the VC dimension to control the complexity of the function class under consideration [31, Chapter 2]. The chaining argument specifically addresses the continuous parameter $z_{2}$, and the complexity of its increments is then controlled using the VC dimension technique applied to the function class specified by the parameters $z_{1},t$. To deal with measurability issues, we use the calculus of outer integral [31, Chapter 1.2]. A crucial part of our proof demonstrates that our process becomes measurable after symmetrization, building on ideas in the monograph [31, Chapter 2.3].
@@ A.1.1 Chaining Argument
Let $\epsilon_{k}=r2^{-k}$ for every $k\geq 0$. An $\epsilon_{k}$-cover of $\mathcal{Z}_{2}$, denoted by $\mathcal{Z}_{2,\epsilon_{k}}$, enables the approximation of every point $z_{2}\in\mathcal{Z}_{2}$ by a point $\pi_{k}(z_{2})\in\mathcal{Z}_{2,\epsilon_{k}}$ such that $\left\|{z_{2}-\pi_{k}(z_{2})}\right\|\leq\epsilon_{k}$. Specifically, we take the initial cover $\mathcal{Z}_{2,\epsilon_{0}}=\{z_{2,0}\}$ to consist solely of the center $z_{2,0}$ of the ball $\mathcal{Z}_{2}$. By employing a standard volume-type argument, we can further select the cover $\mathcal{Z}_{2,\epsilon_{k}}$ such that $\log|\mathcal{Z}_{2,\epsilon_{k}}|\leq p\log(2+r/\epsilon_{k})$ for every $k\geq 1$ [44, Corollary 4.2.13].
@@ A.1.2 Bounding Individual Term using VC-Dimension (Proof of Lemma 3)
In the remainder, we prove Lemma 5. Our proof exploits the fact that the VC dimension of the set family $\mathcal{W}$ controls the Rademacher complexity associated with the indicators $\psi_{z_{1},t}(\xi)$. Additionally, we utilize calculus rules for outer probability measures [31, Section 1.2-5] to address measurability issues. Recall the notion of outer integral [31, Section 1.2].
@@
Let $V_{1/2}$ denote a $1/2$ cover of unit ball in $\mathbb{R}^{d}$ with $\log|V_{1/2}|\leq d\log 6$. Following a standard argument (see, e.g., [44, Exercise 4.2.2]), we know that for every $\bar{v}\in\mathbb{R}^{d}$:
@@
Applying symmetrization techniques from empirical process theory for the outer integral [31, Lemma 2.3.1], we deduce an upper bound for $U(\tau)$:
@@
where $\mathbb{E}^{*}$ is understood as the outer integral on the product probability space involving the data samples $\xi_{1:m}$ and the independent Rademacher random variables $\varepsilon_{1:m}$. To be more formal, we follow [31, Chapter 2.3]. Recall that a single random variable $\xi$ can be defined on the probability space $(\Xi,\mathscr{G},\mathbb{P})$. Thus, we can define our data samples $\xi_{1:m}$ on the product probability space $(\Xi^{\otimes m},\sigma(\mathscr{G}^{\otimes m}),\mathbb{P}^{\otimes m})$, where $\Xi^{\otimes m}$ represents the $m$-fold Cartesian product of $\Xi$, $\sigma(\mathscr{G}^{\otimes m})$ represents the $\sigma$-field generated by the $m$-fold product of $\mathscr{G}$, and $\mathbb{P}^{\otimes m}$ is the $m$-fold product of $\mathbb{P}$, where $\xi_{1},\xi_{2},\ldots,\xi_{m}$ corresponds to the coordinate projections. Then, we introduce Rademacher random variables $\varepsilon_{1:m}=(\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{m})$ defined on another probability space $(\Xi^{\prime},\mathscr{H},\mathbb{P}^{\prime})$. We finally combine these to define the final product probability space of interest:
@@
Lemma 6 is the critical component that resolves the measurability issue. From the empirical process theory, we know that measurability of certain forms is needed at this point, as Fubini’s theorem is not valid for outer integrals [31, Chapter 2.3]. Lemma 6 follows the established ideas to resolve the measurability issues, see, e.g., discussions in the monograph on the measurable class [31, Definition 2.3.3]. The proof of Lemma 6 is deferred to Section A.1.4.
@@
$\mathcal{W}:=\left\{\{\xi\in\Xi:o(z_{1},\xi)\geq t\}\mid z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}\right\}$
@@
For a reference of Sauer-Shelah Lemma, see, e.g., [46, 47] or [44, Theorem 8.3.16].
@@
Since $u_{v}(\xi)$ is $\bar{\sigma}$-sub-exponential, then $\varepsilon_{i}u_{v}(\xi_{i})$ is also $\bar{\sigma}$-sub-exponential for every $v$ by definition. Using [44, Proposition 2.7.1(e)], this implies that, for every $\theta_{2}$, $v$, $\tau\in(0,cm/\bar{\sigma})$:
@@ A.1.4 A Result on Measurability
Recall that $\psi_{z_{1},t}(\xi)=\mathbf{1}\{o(z_{1},\xi)\geq t\}$. First, we note an important observation, which uses the fact that $t\mapsto\mathbf{1}\{w\geq t\}$ is left continuous for every real value $w$. For every $z_{1}$, there exists $t^{o}\in\mathbb{R}^{o}$ with $t^{o}<t$ such that for every $i=1,2,\ldots,m$:
$\psi_{z_{1},t}(\xi_{i})=\psi_{z_{1},t^{o}}(\xi_{i})~{}~{}\text{and}~{}~{}o(z_{1},\xi_{i})-t^{o}\neq 0.$
Fix this $t^{o}$. Next, since $z_{1}\mapsto o(z_{1},\xi)$ is continuous for every $\xi$, this implies the existence of $z_{1}^{o}\in\mathcal{Z}_{1}^{o}$ close to $z_{1}$ such that for every $i=1,2,\ldots,m$:
@@
Appendix B Proof of Theorem 7
@@
$\mathcal{D}(\mathcal{Z}_{S},\mathcal{Z})\leq C\sqrt[4]{\frac{m}{d}}\left\|{\bar{x}}\right\|.$
We shall build on top of these existing results to establish Theorem 7. First, Proposition 13 implies there exists a numerical constant $C_{1}>0$ such that:
@@
Corollary 2 refines Theorem 6 using a standard peeling argument to substitute the norm $\left\|{x}\right\|$ for the constant radius $r$ in the original probability expression in equation (5.2). This adjustment allows varying distances $x$ from the origin, ranging from $\underline{r}$ to $r$, enhancing the applicability of the probability bounds over different scales. For completeness, we give its proof in Section B.1.
@@
We are now ready to prove Theorem 7. We pick
@@
$\begin{split}&\mathbb{P}_{*}\bigg{(}\mathbb{H}(\partial\Phi(x),\partial\Phi_{S}(x))\leq C_{2}\left\|{x}\right\|\cdot\max\{\Delta_{0},\Delta_{0}^{2}\},\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\forall x~{}\text{satisfying}~{}C_{1}\sqrt{d/m}\left\|{\bar{x}}\right\|\leq\left\|{x}\right\|\leq C_{1}\left\|{\bar{x}}\right\|\bigg{)}\geq 1-e^{-d}.\end{split}$ (B.4)
@@
We without loss of generality assume that $C_{2}\max\{\Delta_{0},\Delta_{0}^{2}\}\leq\gamma$ where $\gamma$ is the numerical constant specified in Theorem 12, since otherwise Theorem 7 trivially follows from equation (B.1).
@@
Since we have assumed $C_{2}\max\{\Delta_{0},\Delta_{0}^{2}\}\leq\gamma$, by Theorem 12 we obtain that $x^{*}$ must obey
${\rm dist}(x^{*},\mathcal{Z})\leq C_{2}\left\|{x^{*}}\right\|\cdot\max\{\Delta_{0},\Delta_{0}^{2}\}\leq C_{1}C_{2}\left\|{\bar{x}}\right\|\cdot\max\{\Delta_{0},\Delta_{0}^{2}\}.$
where we recall that $\left\|{x}\right\|\leq C_{1}\left\|{\bar{x}}\right\|$ for every $x\in\mathcal{Z}_{S}$ on this event.
@@
where $C>0$ is an absolute constant. This completes the proof of Theorem 7.
@@ Appendix C Proof of Corollary 1
    where $e(X,\xi)=(h^{sm})^{\prime}(c(X;\xi))\nabla c(X;\xi)$. For every matrix $V\in\mathbb{R}^{D\times D}$, $\langle A,V\rangle$ is $\left\|{V}\right\|_{F}$-subgaussian. Given the property that the sum of two subgaussian random variables is still subgaussian [44, Exercise 2.5.7], we deduce that $\langle A+A^{T},V\rangle$ is $C\left\|{V}\right\|_{F}$-subgaussian for every matrix $V$, where $C>0$ is a universal constant. As a consequence, we derive that, each of
@@
    must be $C\left\|{X_{1}-X_{2}}\right\|_{F}$ subgaussian, and thus $C^{\prime}\left\|{X_{1}-X_{2}}\right\|_{F}$ subexponential for every matrix $V\in\mathbb{R}^{D\times D}$ with $\left\|{V}\right\|_{F}=1$ for some universal constant $C^{\prime}>0$. Here we use the fact that every sub-gaussian distribution is subexponential [44, Section 2.7].
@@
    where $e(X,\xi)=(\psi^{sm})^{\prime}(c(X;\xi))\nabla c(X;\xi)$. Since $A$ is $\sigma$-subgaussian, $\mathbb{E}[\left\|{A}\right\|_{F}]<\infty$. As a result, Assumption C.3 is satisfied with $L(\xi)=2\left\|{A}\right\|_{F}$.
  • (Integrability). Notably $\mathbb{E}[|c(X;\xi)|^{2}]=\mathbb{E}[|b-\langle a,X\rangle|^{2}]<\infty$ for every $x\in\mathbb{R}^{d}$.

### v4

@@ A.1 Proof of Theorem 9
Our proof of Theorem 9 combines two well-established techniques for proving uniform bounds on the supremum of a random process in the statistical learning theory: chaining, which constructs a sequence of progressively finer approximations to the continuous process, and Sauer-Shelah Lemma, a combinatorial principle using the VC dimension to control the complexity of the function class under consideration [34, Chapter 2]. The chaining argument specifically addresses the continuous parameter $z_{2}$, and the complexity of its increments is then controlled using the VC dimension technique applied to the function class specified by the parameters $z_{1},t$. To deal with measurability issues, we use the calculus of outer integral [34, Chapter 1.2]. A crucial part of our proof demonstrates that our process becomes measurable after symmetrization, building on ideas in the monograph [34, Chapter 2.3].
@@ A.1.1 Chaining Argument
Let $\epsilon_{k}=r2^{-k}$ for every $k\geq 0$. An $\epsilon_{k}$-cover of $\mathcal{Z}_{2}$, denoted by $\mathcal{Z}_{2,\epsilon_{k}}$, enables the approximation of every point $z_{2}\in\mathcal{Z}_{2}$ by a point $\pi_{k}(z_{2})\in\mathcal{Z}_{2,\epsilon_{k}}$ such that $\left\|{z_{2}-\pi_{k}(z_{2})}\right\|\leq\epsilon_{k}$. Specifically, we take the initial cover $\mathcal{Z}_{2,\epsilon_{0}}=\{z_{2,0}\}$ to consist solely of the center $z_{2,0}$ of the ball $\mathcal{Z}_{2}$. By employing a standard volume-type argument, we can further select the cover $\mathcal{Z}_{2,\epsilon_{k}}$ such that $\log|\mathcal{Z}_{2,\epsilon_{k}}|\leq p\log(2+r/\epsilon_{k})$ for every $k\geq 1$ [46, Corollary 4.2.13].
@@ A.1.2 Bounding Individual Term using VC-Dimension (Proof of Lemma 3)
In the remainder, we prove Lemma 5. Our proof exploits the fact that the VC dimension of the set family $\mathcal{W}$ controls the Rademacher complexity associated with the indicators $\psi_{z_{1},t}(\xi)$. Additionally, we utilize calculus rules for outer probability measures [34, Section 1.2-5] to address measurability issues. Recall the notion of outer integral [34, Section 1.2].
@@
Let $V_{1/2}$ denote a $1/2$ cover of unit ball in $\mathbb{R}^{d}$ with $\log|V_{1/2}|\leq d\log 6$. Following a standard argument (see, e.g., [46, Exercise 4.2.2]), we know that for every $\bar{v}\in\mathbb{R}^{d}$:
@@
Applying symmetrization techniques from empirical process theory for the outer integral [34, Lemma 2.3.1], we deduce an upper bound for $U(\tau)$:
@@
where $\mathbb{E}^{*}$ is understood as the outer integral on the product probability space involving the data samples $\xi_{1:m}$ and the independent Rademacher random variables $\varepsilon_{1:m}$. To be more formal, we follow [34, Chapter 2.3]. Recall that a single random variable $\xi$ can be defined on the probability space $(\Xi,\mathscr{G},\mathbb{P})$. Thus, we can define our data samples $\xi_{1:m}$ on the product probability space $(\Xi^{\otimes m},\sigma(\mathscr{G}^{\otimes m}),\mathbb{P}^{\otimes m})$, where $\Xi^{\otimes m}$ represents the $m$-fold Cartesian product of $\Xi$, $\sigma(\mathscr{G}^{\otimes m})$ represents the $\sigma$-field generated by the $m$-fold product of $\mathscr{G}$, and $\mathbb{P}^{\otimes m}$ is the $m$-fold product of $\mathbb{P}$, where $\xi_{1},\xi_{2},\ldots,\xi_{m}$ corresponds to the coordinate projections. Then, we introduce Rademacher random variables $\varepsilon_{1:m}=(\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{m})$ defined on another probability space $(\Xi^{\prime},\mathscr{H},\mathbb{P}^{\prime})$. We finally combine these to define the final product probability space of interest:
@@
Lemma 6 is the critical component that resolves the measurability issue. From the empirical process theory, we know that measurability of certain forms is needed at this point, as Fubini’s theorem is not valid for outer integrals [34, Chapter 2.3]. Lemma 6 follows the established ideas to resolve the measurability issues, see, e.g., discussions in the monograph on the measurable class [34, Definition 2.3.3]. The proof of Lemma 6 is deferred to Section A.1.4.
@@
$\mathcal{W}:=\left\{\{\xi\in\Xi:q(z_{1},\xi)\geq t\}\mid z_{1}\in\mathcal{Z}_{1},t\in\mathbb{R}\right\}$
@@
For a reference of Sauer-Shelah Lemma, see, e.g., [48, 49] or [46, Theorem 8.3.16].
@@
Since $u_{v}(\xi)$ is $\bar{\sigma}$-sub-exponential, then $\varepsilon_{i}u_{v}(\xi_{i})$ is also $\bar{\sigma}$-sub-exponential for every $v$ by definition. Using [46, Proposition 2.7.1(e)], this implies that, for every $\theta_{2}$, $v$, $\tau\in(0,cm/\bar{\sigma})$:
@@ A.1.4 A Result on Measurability
Recall that $\psi_{z_{1},t}(\xi)=\mathbf{1}\{q(z_{1},\xi)\geq t\}$. First, we note an important observation, which uses the fact that $t\mapsto\mathbf{1}\{w\geq t\}$ is left continuous for every real value $w$. For every $z_{1}$, there exists $t^{o}\in\mathbb{R}^{o}$ with $t^{o}<t$ such that for every $i=1,2,\ldots,m$:
$\psi_{z_{1},t}(\xi_{i})=\psi_{z_{1},t^{o}}(\xi_{i})~{}~{}\text{and}~{}~{}q(z_{1},\xi_{i})-t^{o}\neq 0.$
Fix this $t^{o}$. Next, since $z_{1}\mapsto q(z_{1},\xi)$ is continuous for every $\xi$, this implies the existence of $z_{1}^{o}\in\mathcal{Z}_{1}^{o}$ close to $z_{1}$ such that for every $i=1,2,\ldots,m$:
@@
Appendix B Proof of Proposition 8
@@
$\mathbb{D}(\mathcal{Z}_{S},\mathcal{Z})\leq C\sqrt[4]{\frac{m}{d}}\left\|{\bar{x}}\right\|.$
We shall build on top of these existing results to establish Proposition 8. First, Proposition 10 and Proposition 12 imply that there exists a numerical constant $C_{1}>0$ such that:
@@
Corollary 4 refines Corollary 1 using a standard peeling argument to substitute the norm $\left\|{x}\right\|$ for the constant radius $r$ in the original probability expression in equation (6.4). This adjustment allows varying distances $x$ from the origin, ranging from $\underline{r}$ to $r$, thus giving bounds on Hausdorff distances between $\partial\Phi$ and $\partial\Phi_{S}$ over different scales. For completeness, we give its proof in Section B.1.
@@
We are now ready to prove Proposition 8. We pick
@@
$\begin{split}&\mathbb{P}_{*}\bigg{(}\mathbb{H}(\partial\Phi(x),\partial\Phi_{S}(x))\leq C_{2}\left\|{x}\right\|\cdot\max\{\Delta_{0},\Delta_{0}^{2}\},\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\forall x~{}\text{satisfying}~{}C_{1}\sqrt{d/m}\left\|{\bar{x}}\right\|\leq\left\|{x}\right\|\leq C_{1}\left\|{\bar{x}}\right\|\bigg{)}\geq 1-e^{-d}\end{split}$ (B.4)
@@
We can without loss of generality assume that $C_{2}\max\{\Delta_{0},\Delta_{0}^{2}\}\leq\gamma$ where $\gamma$ is the numerical constant specified in Theorem 11, since otherwise Proposition 8 trivially follows equation (B.1).
@@
Since we have assumed that $C_{2}\max\{\Delta_{0},\Delta_{0}^{2}\}\leq\gamma$, by Theorem 11 we obtain that $x^{*}$ must satisfy ${\rm dist}(x^{*},\mathcal{Z})\leq C_{3}{\rm dist}(0,\partial\Psi(x^{*}))$ for some numerical constant $C_{3}>0$, and thereby,
$\begin{split}{\rm dist}(x^{*},\mathcal{Z})&\leq C_{3}{\rm dist}(0,\partial\Psi(x^{*}))\leq C_{2}C_{3}\left\|{x^{*}}\right\|\cdot\max\{\Delta_{0},\Delta_{0}^{2}\}\leq C_{1}C_{2}C_{3}\left\|{\bar{x}}\right\|\cdot\max\{\Delta_{0},\Delta_{0}^{2}\}\end{split}$
where the last inequality uses the fact that $\left\|{x}\right\|\leq C_{1}\left\|{\bar{x}}\right\|$ for every $x\in\mathcal{Z}_{S}$ on this event.
@@
where $C>0$ is an absolute constant. This completes the proof of Proposition 8.
@@ Appendix C Proof of Corollary 2
  • (Assumption C.1). By definition,
    $\nabla c(x,\xi)=\begin{bmatrix}\partial_{y}c(x,\xi)\\ \partial_{w}c(x,\xi)\end{bmatrix}=\begin{bmatrix}\langle v,w\rangle u\\ \langle u,y\rangle v\end{bmatrix},~{}~{}~{}(h^{sm})^{\prime}(c(x;\xi))\nabla c(x;\xi)=-\nabla c(x,\xi).$
    At the center $x=(0,0)\in\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}$, the gradients vanish $\nabla c(0,\xi)=(h^{sm})^{\prime}(c(0;\xi))\nabla c(0;\xi)=0$.
  • (Assumption C.2). It is easy to verify that, for $x_{1}=(y_{1},w_{1})$ and $x_{2}=(y_{2},w_{2})$:
    $\begin{split}\nabla c(x_{1},\xi)-\nabla c(x_{2},\xi)&=\begin{bmatrix}\langle v,w_{1}-w_{2}\rangle u\\ \langle u,y_{1}-y_{2}\rangle v\end{bmatrix}\\ e(x_{1},\xi)-e(x_{2},\xi)&=\begin{bmatrix}\langle v,w_{2}-w_{1}\rangle u\\ \langle u,y_{2}-y_{1}\rangle v\end{bmatrix}\end{split},$
    where $e(x,\xi)=(h^{sm})^{\prime}(c(x;\xi))\nabla c(x;\xi)$. Given the property that the product of two subgaussian random variables is subexponential [46, Lemma 2.7.7], and that the sum of two subexponential random vectors remain subexponential [46, Lemma 2.7.11], one can easily show that $\nabla c(x_{1},\xi)-\nabla c(x_{2},\xi)$ and $e(x_{1},\xi)-e(x_{2},\xi)$ are $C\sigma^{2}\left\|{x_{1}-x_{2}}\right\|$ subexponential vectors where $C$ is absolute.
  • (Assumption C.3). It is easy to verify that
    $\begin{split}\left\|{\nabla c(x_{1},\xi)-\nabla c(x_{2},\xi)}\right\|&\leq 2\left\|{u}\right\|\left\|{v}\right\|\left\|{x_{1}-x_{2}}\right\|\\ \left\|{e(x_{1},\xi)-e(x_{2},\xi)}\right\|&\leq 2\left\|{u}\right\|\left\|{v}\right\|\left\|{x_{1}-x_{2}}\right\|\end{split},$
    where $e(x,\xi)=(h^{sm})^{\prime}(c(x;\xi))\nabla c(x;\xi)$. Since $u,v$ are independent $\sigma$-subgaussian random vectors, $\mathbb{E}[L(\xi)]=2\mathbb{E}[\left\|{u}\right\|\left\|{v}\right\|]<\infty$. Thus, Assumption C.3 is satisfied with $L(\xi)=2\left\|{u}\right\|\left\|{v}\right\|$.
  • (Integrability). Notably $\mathbb{E}[h(c(x;\xi))]=\mathbb{E}[|b-\langle u,y\rangle\langle v,w\rangle|]<\infty$ for every $x=(y,w)\in\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}$.
=$\mathcal{F}=\left\{\{u\in\mathbb{R}^{d_{1}},v\in\mathbb{R}^{d_{2}},b\in\mathbb{R}:\langle u,y\rangle\langle v,w\rangle-b\geq t\}\mid y\in\mathbb{R}^{d_{1}},w\in\mathbb{R}^{d_{2}},t\in\mathbb{R}\right\}.$
We now bound its VC dimension using Theorem 6. Note $(y,w)\mapsto\langle u,y\rangle\langle v,w\rangle-b$ is a degree $2$ polynomial in $(y,w)$ for every $(u,v,b)$. Theorem 6 implies for some universal constant $C>0$:
${\rm vc}(\mathcal{F})\leq C(d_{1}+d_{2})\log(d_{1}+d_{2}).$
Corollary 2 then follows by applying Theorem 5.
Appendix D Proof of Corollary 3
@@
    where $e(X,\xi)=(h^{sm})^{\prime}(c(X;\xi))\nabla c(X;\xi)$. For every matrix $V\in\mathbb{R}^{D\times D}$, $\langle A,V\rangle$ is $\sigma\left\|{V}\right\|_{F}$-subgaussian. Given the property that the sum of two subgaussian random variables is still subgaussian [46, Exercise 2.5.7], we deduce that $\langle A+A^{T},V\rangle$ is $C\left\|{V}\right\|_{F}$-subgaussian for every matrix $V$, where $C>0$ is a universal constant. As a consequence, we derive that, each of
@@
    must be $C\sigma\left\|{X_{1}-X_{2}}\right\|_{F}$ subgaussian, and thus $C^{\prime}\sigma\left\|{X_{1}-X_{2}}\right\|_{F}$ subexponential for every matrix $V\in\mathbb{R}^{D\times D}$ with $\left\|{V}\right\|_{F}=1$ for some universal constant $C^{\prime}>0$. Here we use the fact that every sub-gaussian distribution is subexponential [46, Section 2.7].
@@
    where $e(X,\xi)=(h^{sm})^{\prime}(c(X;\xi))\nabla c(X;\xi)$. Since $A$ is $\sigma$-subgaussian, $\mathbb{E}[\left\|{A}\right\|_{F}]<\infty$. As a result, Assumption C.3 is satisfied with $L(\xi)=2\left\|{A}\right\|_{F}$.
  • (Integrability). Notably $\mathbb{E}[h(c(X;\xi))]=\mathbb{E}[|b-\langle A,X\rangle^{2}|]<\infty$ for every $x\in\mathbb{R}^{d}$.